Hello!
I would like to know the formula (and answer) for the following:
I trade currency on a daily basis (the market is open for 5 days per 7 day week), so if I began with $200, and I gained 10% on my principal on a DAILY basis, how long would it take me to reach $10,000?
For example, Monday = $200, Tuesday I would have $220 (200*10%=220), Wednesday I would have $242 (220*10%=242), etc. Is there a simple formula for determining how long it would take me to achieve a $10,000 goal, assuming I can make 10% gains every trading day using the previous days total as my new principal.
Thank you!

Answer by solve_for_x(190)

The formula that you are looking for is:

FV = PV * (1 + 0.10)^n

where FV is the future value ($10,000), PV is the present value ($200), and n is the number of
trading days.

10,000 = 200(1.10)^n

(1.10)^n = 10000/200

(1.10)^n = 50

Taking the natural logarithm of each side gives you:

n * ln(1.10) = ln(50)

n = ln(50) / ln(1.10)

n = 41.05 trading days

Note: If your calculator does not have the natural logarithm
function (ln), you can use the regular base-10 logarithm
function (log). As long as you use the same function in both
places, you'll get the same answer.